# Identification of Dynamic Nonlinear Panel Models

### v2

@@ 5 Censored Dependent Variable Model
where $X_{it}$ represents the endogenous covariate, $Z_{it}$ stands for the exogenous covariate that satisfies Assumption 1, $Y^{*}_{it}$ denotes the latent outcome not observed in the data, and $Y_{it}$ represents the observed outcome, which is censored at zero. The threshold for censoring can be replaced with other nonzero constants. The identification strategy is still to derive conditions for $\theta_{0}$ based on the stationarity assumption. Here the conditional distribution of $\alpha_{i}+\epsilon_{it}\mid Z_{st}$ is not identified since we only observe the censoring outcome $Y_{it}$. While we can still bound the conditional distribution $\alpha_{i}+\epsilon_{it}\mid Z_{st}$ using variation in the joint distribution $(Y_{it},Z_{it},X_{it})$. The following proposition presents the identification results of $\theta_{0}$ with censoring outcomes.
@@
Under Assumption 1, an identified set for $\theta_{0}$ is the set of parameters $\theta$ that satisfy the following condition:
$\Pr(0<Y_{is}\leq z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma-a\mid z_{st})+\Pr(Y_{is}=0\mid z_{st})\geq\Pr(Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-a\mid z_{st}),$
for any $a\in\mathcal{R}$, any $(s,t)\leq T$, and any $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
@@
Similar to discrete choice models studied in previous sections, Proposition 8 characterizes an identified set for $\theta_{0}$ by exploiting the variation in the joint distribution $(Y_{it},X_{it})\mid Z_{ist}$ over time and the variation in the exogenous covariates $Z_{ist}$. The bounds on the distribution $\alpha_{i}+\epsilon_{it}\mid Z_{st}$ can be derived either from the probability $\Pr(0<Y_{it}\leq y\mid z_{st})$ or $\Pr(Y_{it}=0\mid z_{st})$, depending on the value of the covariate index $z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma$. This result still accommodates both static and dynamic models with censored outcomes.
@@
Static model: consider that the standard stationarity assumption holds, where $\epsilon_{it}\mid\alpha_{i},W_{ist}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{is}\mid\alpha_{i},W_{ist}$. Then, the identifying condition in Proposition 8 is given as
$\Pr(0<Y_{is}\leq w_{s}^{\prime}\theta-a)+\Pr(Y_{is}=0\mid w_{st})\geq\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}).$ (11)
@@
The above restriction is informative only when $w_{t}^{\prime}\theta\geq a$, otherwise the lower bound becomes zero. We discuss two cases for the constant $a$. When $w_{s}^{\prime}\theta\leq a\leq w_{t}^{\prime}\theta$, then condition (11) becomes
$\Pr(Y_{is}=0\mid w_{st})\geq\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}).$
When $a$ satisfies $a\leq\min\{w_{s}^{\prime}\theta,w_{t}^{\prime}\theta\}$, condition (11) transforms into
$\Pr(0<Y_{is}\leq w_{s}^{\prime}\theta-a\mid w_{st})+\Pr(Y_{is}=0\mid w_{st})=\Pr(Y_{is}\leq w_{s}^{\prime}\theta-a\mid w_{st})\\ \geq\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}).$
@@
Since the above condition needs to hold for any $(s,t)$ and is symmetric in $(s,t)$, it becomes equalities after exchanging $s$ and $t$.
The following lemma derives the identified set for $\theta_{0}$ using both conditional moment inequalities and equalities.
Lemma 2.
Assuming that $\epsilon_{it}\mid\alpha_{i},W_{ist}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{is}\mid\alpha_{i},W_{ist}$, an identified set for $\theta_{0}$ is characterized by the set of parameters $\theta$ that satisfy the following conditions:
$&\text{If }w_{s}^{\prime}\theta\leq a\leq w_{t}^{\prime}\theta\Longrightarrow\Pr(Y_{is}=0\mid w_{st})\geq\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}),\\ &\text{If }a\leq\min\{w_{s}^{\prime}\theta,w_{t}^{\prime}\theta\}\Longrightarrow\Pr(Y_{is}\leq w_{s}^{\prime}\theta-a\mid w_{st})=\Pr(Y_{it}\leq w_{t}^{\prime}\theta-a\mid w_{st}),\end{aligned}\right.$
for any $s,t\leq T$ and any $w_{st}\in\mathcal{R}^{d}\times\mathcal{R}^{d}$.
@@
6 Generalization to Nonseparable Models
The key idea underlying our identification strategy generalizes further beyond the models considered before. In this section, we show that our approach relies only on two conditions: partial stationarity and “index monotonicity.” Additional model specifications and restrictions, such as the scalar-additivity of fixed effects and time-varying errors, are not needed for our identification strategy.
$Y_{it}$ $=G\left(W_{it}^{{}^{\prime}}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right),$
where $\alpha_{i}$ is the individual fixed effect of arbitrary dimension, $\epsilon_{it}$ is the time-varying error of arbitrary dimension, $W_{it}$ is a vector of observable covariates, $\theta_{0}\in\mathcal{R}^{d}$ is a conformable vector of parameter, and the function $G$ is allowed to be unknown, nonseparable but assumed to be weakly monotone in the the parametric index:
Assumption 4 (Monotonicity).
Note that, by setting $\alpha_{i},\epsilon_{it}$ to be scalar-valued, and $G\left(W_{it}^{{}^{\prime}}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)=\mathbf{\mathbbm{1}}\left\{W_{it}^{{}^{\prime}}\theta_{0}+\alpha_{i}+\epsilon_{it}\geq 0\right\}$, we obtain the dynamic binary choice model in Section 2, where $G$ is by construction weakly increasing in $W_{it}^{{}^{\prime}}\theta_{0}$.
As before, we decompose $W_{it}$, and correspondingly $\theta_{0}$, into two components, $W_{it}=\left(Z_{it}^{{}^{\prime}},X_{it}^{{}^{\prime}}\right)^{{}^{\prime}}$ and $\theta_{0}=\left(\beta_{0}^{{}^{\prime}},\gamma_{0}^{{}^{\prime}}\right)^{{}^{\prime}}$so that
where $Z_{it}$ denotes exogenous covariates while $X_{it}$ denotes endogenous covariates, with the precise definition of exogeneity encoded by the partial stationarity in Assumption 1.
Now we show how partial stationarity can be exploited in conjunction with weak monotonicity (Assumption 4) to obtain identifying restrictions in the presence of endogeneity.
Let $\mathcal{Y}$ denote the support of $Y_{it}$. For any $a\in\mathcal{R}$ and $y\in\mathcal{Y}$, conditional on $W_{ist}=w_{st}$, we can bound $\mathbf{\mathbbm{1}}\left\{G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\}$ below as
$\mathbf{\mathbbm{1}}\left\{G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\}$ $\geq\mathbf{\mathbbm{1}}\left\{w_{t}^{\prime}\theta_{0}\leq a\right\}\mathbf{\mathbbm{1}}\left\{G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\}$
$\geq\mathbf{\mathbbm{1}}\left\{w_{t}^{\prime}\theta_{0}\leq a\right\}\mathbf{\mathbbm{1}}\left\{G\left(w_{t}^{\prime}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\},$
and above as
$\mathbf{\mathbbm{1}}\left\{G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\right\}$
[…]
=\ = $1-\mathbf{\mathbbm{1}}\left\{w_{t}^{\prime}\theta_{0}>a\right\}\mathbf{\mathbbm{1}}\left\{G\left(w_{t}^{\prime}\theta_{0},\,\alpha_{i},\,\epsilon_{it}\right)<y\right\}.$
Then, the conditional probability $\mathbb{P}\left(G\left(a,\,\alpha_{i},\,\epsilon_{it}\right)\geq y\mid w_{st}\right)$ given $W_{ist}=w_{st}$ can be bounded as follows:
[…]
Taking conditional expectation of the above with respect to $X_{ist}$ given $Z_{ist}=z_{st}$ yields
$\mathbb{P}\left(Y_{is}\geq y,W_{is}^{\prime}\theta_{0}\leq a\mid z_{st}\right)$
[…]
The key difference of (6) relative to the corresponding identifying restrictions in previous sections lies in that the “middle term” in (6) is no longer the conditional CDF of $\alpha_{i}+\epsilon_{it}$, but a conditional probability about the event $G\left(a,\,\alpha_{i},\,\epsilon_{is}\right)\geq y$, with the latter representation not dependent on scalar-additivity of fixed effects and time-varying errors.
Proposition 9.
Under Assumptions 1 and 4, the following condition holds:
$1-\mathbb{P}\left(Y_{it}<y,z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}>a\mid z_{st}\right)\geq\mathbb{P}\left(Y_{is}\geq y,z_{s}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}\leq a\mid z_{st}\right),$ (12)
for any $a\in\mathcal{R},$ $y\in\mathcal{Y}$, $s\neq t\leq T$, and $z_{st}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{z}}$.
Note that in the binary choice setting of Section 2, it suffices to set $y=1$ in (12), which then coincides with the identifying condition (7) in Proposition 1. It is also straightforward to show that our identification results in Lemma 9 can be adapted to the ordered response model in Section 3, the multinomial choice model in Section 4, and the censored outcome model in Section 5 without the scalar additive specifications.
@@ 7 Simulation
This section examines the finite sample performance of our identification approaches using Monte Carlo simulations. Since the literature has extensively studied binary choice models, we focus on the static and dynamic ordered choice models explored in Section 3 as examples to illustrate the approach. We implement the kernel-based CLR inference approach proposed in the papers by Chernozhukov, Lee, and Rosen (2013) and Chen and Lee (2019), developed to construct confidence interval based on general conditional moment inequalities.
@@ 7.1 Static Ordered Choice Model
We consider the covariates $Z_{it}$ to be exogenous and Lemma 1 characterizes the identified set for $\beta_{0}$ using the following conditional moment inequalities:
$E[g(Z_{ist},Y_{ist};\beta_{0})\mid z_{st}]\geq 0,$
@@
$g(Z_{ist},Y_{ist};\beta_{0})=\left\{\begin{aligned} &\mathbf{\mathbbm{1}}\{b_{2}-Z_{is}^{\prime}\beta\geq b_{2}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}=1\}-\mathbf{\mathbbm{1}}\{Y_{it}=1\});\\ &\mathbf{\mathbbm{1}}\{b_{2}-Z_{is}^{\prime}\beta\geq b_{3}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}=1\}-\mathbf{\mathbbm{1}}\{Y_{it}\in\{1,2\}\});\\ &\mathbf{\mathbbm{1}}\{b_{3}-Z_{is}^{\prime}\beta\geq b_{2}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}\in\{1,2\}\}-\mathbf{\mathbbm{1}}\{Y_{it}=1\});\\ &\mathbf{\mathbbm{1}}\{b_{3}-Z_{is}^{\prime}\beta\geq b_{3}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}\in\{1,2\}\}-\mathbf{\mathbbm{1}}\{Y_{it}\in\{1,2\}\}).\end{aligned}\right.$
@@ 7.2 Dynamic Ordered Choice Model
for $1\leq t\leq T$. The initial value $Y_{i0}\in\{1,2,3\}$ is generated independently of all variables, and follow the distribution $\Pr(Y_{i0}=1)=0.6,\Pr(Y_{i0}=2)=\Pr(Y_{i0}=3)=0.2$.
@@
    $1-\sum_{j=2}^{3}\Pr\left(Y_{i1}=y_{j}\mid z,y_{0}\right)*\mathbf{\mathbbm{1}}$$\{b_{j}-z_{1}^{\prime}\beta-y_{0}\gamma\geq a\}$
    $\geq\sum_{j=1}^{2}\Pr\left(Y_{is}=y_{j},b_{j+1}-z_{s}^{\prime}\beta-Y_{is-1}\gamma\leq a\mid z,y_{0}\right),$
    $1-\sum_{j=2}^{3}\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma$ $\geq a\mid z,y_{0})$
    $\geq\sum_{j=1}^{2}\Pr$ $\left(Y_{i1}=y_{j}\mid z,y_{0}\right)*\mathbf{\mathbbm{1}}\{b_{j+1}-z_{1}^{\prime}\beta-y_{0}\gamma\leq a\},$
    for any $a\in\{b_{j}-z_{1}^{\prime}\beta-y_{0}\gamma,b_{j}-z_{s}^{\prime}\beta-\gamma,b_{j}-z_{s}^{\prime}\beta-2\gamma,b_{j}-z_{s}^{\prime}\beta-3\gamma\}_{j=2}^{T}$;
@@
    $1-\sum_{j=2}^{3}\Pr(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma$ $\geq a\mid z,y_{0})$
    $\geq\sum_{j=1}^{2}$ $\Pr\left(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-Y_{it-1}\gamma\leq a\mid z,y_{0}\right),$
    for any $a\in\{b_{j}-z_{s}^{\prime}\beta-\gamma,b_{j}-z_{s}^{\prime}\beta-2\gamma,b_{j}-z_{s}^{\prime}\beta-3\gamma,b_{j}-z_{t}^{\prime}\beta-\gamma,b_{j}-z_{t}^{\prime}\beta-2\gamma,b_{j}-z_{t}^{\prime}\beta-3\gamma\}_{j=2}^{3}$.

### v3

@@ 3.4 Censored Outcome Model
where $Y^{*}_{it}$ denotes the latent outcome which is not observed in the data, and $Y_{it}$ represents the observed outcome, censored at zero. The threshold for censoring can be replaced with other nonzero constants.
The identification strategy is still to exploit the partial stationarity assumption and bound the conditional distribution of $\alpha_{i}+\epsilon_{it}\mid Z_{i}=z$. This censored outcome model imposes an additional structure between the outcome and the parametric index: when $Y_{it}>0$, we have $Y_{it}=Y^{*}_{it}$ and
$\alpha_{i}+\epsilon_{it}\leq c\Longleftrightarrow Y_{it}-Z_{it}^{\prime}\beta_{0}-X_{it}^{\prime}\gamma_{0}\leq c.$
This specific structure can be exploited to further tighten the identified set for $\theta_{0}$, and we provide the details of the identification strategy in Appendix A.7. The following proposition presents the identification results of $\theta_{0}$ with censored outcomes.
@@
Under Assumption 1, $\theta_{0}\in\Theta_{I,cen}$, where the identified set $\Theta_{I,cen}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
$\max_{t=1,...,T}\mathbb{P}(Y_{it}\leq z_{t}^{\prime}\beta+X_{it}^{\prime}\gamma-c\mid z)\leq\max_{s=1,...,T}\left\{\mathbb{P}(0<Y_{is}\leq z_{s}^{\prime}\beta+X_{is}^{\prime}\gamma-c\mid z)+\mathbb{P}(Y_{is}=0\mid z)\right\},$
for any $c\in\mathcal{R}$ and any realization $z=(z_{1},...,z_{T})$ in the support of $Z_{i}$.
@@
Similar to discrete choice models studied in previous sections, Proposition 9 characterizes an identified set for $\theta_{0}$ by exploiting the variation in the joint distribution of $(Y_{it},X_{it})\mid Z_{i}$ over time and the variation in the exogenous covariates $Z_{i}$. The bounds on the distribution $\alpha_{i}+\epsilon_{it}\mid Z_{i}=z$ can be derived either from the probability $\mathbb{P}(0<Y_{it}\leq y\mid z)$ or $\mathbb{P}(Y_{it}=0\mid z)$, depending on the value of the covariate index $z_{t}^{\prime}\beta_{0}+X_{it}^{\prime}\gamma_{0}$. This result still accommodates both static and dynamic models with censored outcomes, and we provide the simplified results for each model.
@@
Static model: consider that the full stationarity assumption holds, i.e., $\epsilon_{it}\mid\alpha_{i},W_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{is}\mid\alpha_{i},W_{i}$. Then, the identifying condition in Proposition 9 is given as
$\mathbb{P}(Y_{it}\leq w_{t}^{\prime}\theta-c\mid w)\leq\mathbb{P}(0<Y_{is}\leq w_{s}^{\prime}\theta-c\mid w)+\mathbb{P}(Y_{is}=0\mid w).$ (25)
@@
The above restriction is informative only when $w_{t}^{\prime}\theta-c\geq 0$, otherwise the lower bound becomes zero. We discuss two cases for the constant $c$: $w_{s}^{\prime}\theta\leq c\leq w_{t}^{\prime}\theta$ and $c\leq\min\{w_{s}^{\prime}\theta,w_{t}^{\prime}\theta\}$.
When $w_{s}^{\prime}\theta\leq c\leq w_{t}^{\prime}\theta$, then condition (25) becomes
$\mathbb{P}(Y_{it}\leq w_{t}^{\prime}\theta-c\mid w)\leq\mathbb{P}(Y_{is}=0\mid w).$
When $c$ satisfies $c\leq\min\{w_{s}^{\prime}\theta,w_{t}^{\prime}\theta\}$, condition (25) transforms into
$\mathbb{P}(Y_{it}\leq w_{t}^{\prime}\theta-c\mid w)\leq\mathbb{P}(0<Y_{is}\leq w_{s}^{\prime}\theta-c\mid w)+\mathbb{P}(Y_{is}=0\mid w)=\mathbb{P}(Y_{is}\leq w_{s}^{\prime}\theta-c\mid w).$
@@
Since the above condition needs to hold for any $(s,t)$ and is symmetric in $(s,t)$, it becomes equalities after exchanging $s$ and $t$. The following lemma summarizes the results for the static model.
Corollary 2.
Assuming that $\epsilon_{is}\mid\alpha_{i},W_{i}\stackrel{{\scriptstyle d}}{{\sim}}\epsilon_{it}\mid\alpha_{i},W_{i}$, the identified set $\Theta_{I,cen}$ consists of all $\theta=\left(\beta^{{}^{\prime}},\gamma^{{}^{\prime}}\right)^{{}^{\prime}}\in\mathcal{R}^{d_{z}}\times\mathcal{R}^{d_{x}}$ such that
$&\text{If }w_{s}^{\prime}\theta\leq c\leq w_{t}^{\prime}\theta\Longrightarrow\mathbb{P}(Y_{it}\leq w_{t}^{\prime}\theta-c\mid w)\leq\mathbb{P}(Y_{is}=0\mid w);\\ &\text{If }c\leq\min\{w_{s}^{\prime}\theta,w_{t}^{\prime}\theta\}\Longrightarrow\mathbb{P}(Y_{it}\leq w_{t}^{\prime}\theta-c\mid w)=\mathbb{P}(Y_{is}\leq w_{s}^{\prime}\theta-c\mid w),\end{aligned}\right.$
for any $c\in\mathcal{R}$, any $(s,t)\leq T$, and any realization $w=(w_{1},...,w_{T})$ in the support of $W_{i}$.
@@ 4 Simulation
This section examines the finite sample performance of our identification approaches using Monte Carlo simulations. We focus on the static and dynamic ordered choice models explored in Section 3.2 as examples to illustrate the approach. We implement the kernel-based CLR inference approach proposed in the papers by Chernozhukov, Lee, and Rosen (2013) and Chen and Lee (2019), developed to construct confidence interval based on general conditional moment inequalities.
@@ 4.1 Static Ordered Choice Model
With $Y_{i}:=(Y_{i1},Y_{i2})$ and $Z_{i}:=(Z_{i1},Z_{i2})$, Corollary 1 characterizes the identified set for $\beta_{0}$ using the following conditional moment inequalities: for $s\neq t\leq 2$,
$E[g(Z_{i},Y_{i};\beta_{0})\mid z]\geq 0,$
@@
$g(Z_{i},Y_{i};\beta_{0})=\left\{\begin{aligned} &\mathbf{\mathbbm{1}}\{b_{2}-Z_{is}^{\prime}\beta\geq b_{2}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}=1\}-\mathbf{\mathbbm{1}}\{Y_{it}=1\});\\ &\mathbf{\mathbbm{1}}\{b_{2}-Z_{is}^{\prime}\beta\geq b_{3}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}=1\}-\mathbf{\mathbbm{1}}\{Y_{it}\in\{1,2\}\});\\ &\mathbf{\mathbbm{1}}\{b_{3}-Z_{is}^{\prime}\beta\geq b_{2}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}\in\{1,2\}\}-\mathbf{\mathbbm{1}}\{Y_{it}=1\});\\ &\mathbf{\mathbbm{1}}\{b_{3}-Z_{is}^{\prime}\beta\geq b_{3}-Z_{it}^{\prime}\beta_{0}\}(\mathbf{\mathbbm{1}}\{Y_{is}\in\{1,2\}\}-\mathbf{\mathbbm{1}}\{Y_{it}\in\{1,2\}\}).\end{aligned}\right.$
@@ 4.2 Dynamic Ordered Choice Model
for $1\leq t\leq T$. The initial value $Y_{i0}\in\{1,2,3\}$ is generated independently of all variables, and follow the distribution $\mathbb{P}(Y_{i0}=1)=0.6,\mathbb{P}(Y_{i0}=2)=\mathbb{P}(Y_{i0}=3)=0.2$.
@@
    $1-\sum_{j=2}^{3}\mathbb{P}\left(Y_{i1}=y_{j}\mid z,y_{0}\right)*\mathbf{\mathbbm{1}}$$\{b_{j}-z_{1}^{\prime}\beta-y_{0}\gamma\geq c\}$
    $\geq\sum_{j=1}^{2}\mathbb{P}\left(Y_{is}=y_{j},b_{j+1}-z_{s}^{\prime}\beta-Y_{is-1}\gamma\leq c\mid z,y_{0}\right),$
    $1-\sum_{j=2}^{3}\mathbb{P}(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma$ $\geq c\mid z,y_{0})$
    $\geq\sum_{j=1}^{2}\mathbb{P}$ $\left(Y_{i1}=y_{j}\mid z,y_{0}\right)*\mathbf{\mathbbm{1}}\{b_{j+1}-z_{1}^{\prime}\beta-y_{0}\gamma\leq c\},$
    for any $c\in\{b_{j}-z_{1}^{\prime}\beta-y_{0}\gamma,b_{j}-z_{s}^{\prime}\beta-\gamma,b_{j}-z_{s}^{\prime}\beta-2\gamma,b_{j}-z_{s}^{\prime}\beta-3\gamma\}_{j=2}^{T}$;
@@
    $1-\sum_{j=2}^{3}\mathbb{P}(Y_{is}=y_{j},b_{j}-z_{s}^{\prime}\beta-Y_{is-1}\gamma$ $\geq c\mid z,y_{0})$
    $\geq\sum_{j=1}^{2}$ $\mathbb{P}\left(Y_{it}=y_{j},b_{j+1}-z_{t}^{\prime}\beta-Y_{it-1}\gamma\leq c\mid z,y_{0}\right),$
    for any $c\in\{b_{j}-z_{s}^{\prime}\beta-\gamma,b_{j}-z_{s}^{\prime}\beta-2\gamma,b_{j}-z_{s}^{\prime}\beta-3\gamma,b_{j}-z_{t}^{\prime}\beta-\gamma,b_{j}-z_{t}^{\prime}\beta-2\gamma,b_{j}-z_{t}^{\prime}\beta-3\gamma\}_{j=2}^{3}$.